# EvTTC: An Event Camera Dataset for Time-to-Collision Estimation

### v1

@@ VI TTC Benchmark
STRTTC [24] demonstrates robust performance under various conditions, attributed to its resilient sampling strategy and efficient initialization scheme. In contrast, CMax [32] achieves higher accuracy by utilizing all events within the bounding box, enhancing the signal-to-noise ratio. However, its nonlinear least-squares optimization adds computational complexity, with processing 3e5 events typically taking around 3 seconds. As an event-based incremental methodology, ETTCM [33] does not scale well with input data generated per unit time. FAITH [34] replaces optical flow with normal flow for TTC estimation, leading to worse accuracy. AEB-Tracker [35] achieves exceptionally high computational efficiency, reaching up to 10 kHz. However, under conditions with significant background noise, the algorithm requires a low-frequency detector to identify regions (e.g., two taillights) capable of generating event blobs. Image’s FOE [20] achieves high accuracy and efficiency across all sequences; Nevertheless, its performance is constrained by the limitations of standard cameras (e.g., data latency, motion blur).
@@ Acknowledgment
We thank Javier Hidalgo-Carrió and Davide Scaramuzza for releasing the design of the Beamsplitter [36], based on which we build our FCW system. We thank Mr. Yanggang Sheng and Mr. Junkai Niu for the help in the data collection.

### v3

@@ VI TTC Benchmark
STRTTC [32] demonstrates robust performance under various conditions, attributed to its resilient sampling strategy and efficient initialization scheme. Among event-based methods, CMax [28] achieves the highest accuracy by utilizing all events within the bounding box, thereby enhancing the signal-to-noise ratio. However, its nonlinear least-squares optimization adds computational complexity, with processing 3e5 events typically taking around 3 seconds. ETTCM [30] estimates the TTC and reports computation time per event. For easier comparison, we define its computation time as the product of the time per computation and the total number of events processed. In our experiments, while it shows high efficiency, it does not scale well with input data generated per unit time. Using event-based normal flow as input, FAITH [27] employs a RANSAC-based algorithm to determine the FOE, demonstrating real-time performance. However, substituting normal flow for optical flow leads to a decrease in the accuracy of TTC estimation.
AEB-Tracker [33] achieves exceptionally high computational efficiency, reaching up to 10 kHz. However, under conditions with significant background noise, the algorithm requires a low-frequency detector to identify regions (e.g., two taillights) capable of generating event blobs. Furthermore, it requires that the tracked event blobs adhere to a spatio-temporal Gaussian likelihood model [33]; otherwise, the accuracy of the estimated TTC may decrease. Image’s FoE [20] takes RGB images from a camera running at 20 Hz as input. It achieves high accuracy and efficiency across all sequences. Nevertheless, its performance is constrained by the limitations of standard cameras (e.g., data latency, motion blur).
@@ Acknowledgment
We thank Javier Hidalgo-Carrió and Davide Scaramuzza for releasing the design of the Beamsplitter [41], based on which we build our FCW system. We thank Mr. Yanggang Sheng and Mr. Junkai Niu for the help in the data collection. We also thank Mr. Sheng Zhong for proofreading.
Supplementary Material
The figure below demonstrates detailed test results for the TTC estimation methods described in Sec. VI, evaluated on our dataset.
(a) CCRs1-low
(b) CCRs1-medium
(c) CCRs1-high
(d) CCRs2-low
(e) CCRs2-medium
(f) CCRs2-high
(g) CCRm-low
(h) CCRm-medium
(i) slider-750
(j) slider-1000